Shift report 20010909 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Sep 09 2001 - 15:51:43 EDT

  • Next message: Apache: "Shift report 20010910 16:00-24:00"

    ------------------------------------------------------------
    Date: Sun Sep  9 15:51:43 2001, Shift: 08:00-16:00
    Supervisor: ZM, with: -
    Log book#8 pages 98 - 100
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam since 10:10 with B=18.2x10^9, Y=21.6x10^9
       at the begining. Still runing.
    2.Run #4650 - high ZDC rates (~200) and to high dead  time.
      Problems to write data to HPSS (called to RCF person on 
      shift - problem was solved).
      Run #4652 with triggers: (1(1:2), 3(1:1), 4(1:6) and
      collected 226K events. Now, about 1M events with this
      setting.
      Run #4653 at D4 - off (run for T4 and T5 test).
                100K event collected.
      Run #4654 New setting: FS  4 deg. (magnets 1/5 A)
                            MRS 40 deg. (magnet 3B)  
    3.No detector problems although a few T3 connections
      have to be inspected during next access to IR.
    
    * This is my last shift during this emergency (T3) visit
      to the BNL. Cheers - Zbigniew *
      
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Sep 09 2001 - 15:52:13 EDT